OpenStudy (leozap1):

There are six girls and eight boys on a high school debate team Two students from the team are chosen at random to attend a debate workshop. What is the probability that at least one of the students chosen is a girl?

OpenStudy (leozap1):

How do I solve these types of problems?

OpenStudy (ybarrap):

Let g be the event a girl is chosen and b the event that a boy is chosen To have at least one girl in 2 tries you can have gg bg or gb Where gg means a girl was chosen first and second and bg means a boy was chosen first and a girl chosen second Do yo see that?

OpenStudy (leozap1):

Okay, but how do I find the probability that a girl is selected at least once?

OpenStudy (ybarrap):

On the first try what is the probability that a person chosen is a girl? You have 6 girls and 8 boys, right? What fraction of these are girls?

OpenStudy (leozap1):

There are 14 total people, so there is 6 chances a girl is chosen out of 14. So 6/14?

OpenStudy (ybarrap):

correct! So for the second selection, you will not have 14 people, but 13 because the girl was chosen on the first slot. So now there will be 5 girls and 8 boys right? At this point, what is the chance of choosing a girl?

OpenStudy (leozap1):

5/13 but why would you assume a girl was chosen rather than a boy on the first pick?

OpenStudy (ybarrap):

Because we are saying, what if we choose a girl on the first then on the second. We will take care of the case where a boy was selected first next. Now, what we have done is figured out the chance of gg, which is $$ \cfrac{6}{14}\times\cfrac{5}{13} $$ Let's take the next case, bg. What is the chance that the 1st choice is a boy? We have 6 girls and 8 boys at this point.

OpenStudy (leozap1):

So 6/14 again

OpenStudy (ybarrap):

No, 8/14 - do you see why? This is the chance of a boy being selected first What we are doing is computing mutually exclusive cases. All cases are gg,gb,bg,bb, right? So the probability, P(gg) + P(gb) + P(bg) + P(bb) = 1. Do you see that? We just care about the case where there is at least one girl. that's only 3 of the 4 cases. In fact you can almost see the answer without all these computations. I'll show you that shortcut in a sec. Let's compute P(bg): $$ \cfrac{8}{14}\times\cfrac{6}{13} $$ Do you see why?

OpenStudy (leozap1):

Okay, I think I understand now.

OpenStudy (ybarrap):

Next, what is P(gb)? That's the last one

OpenStudy (leozap1):

It would be 6/14 * 8/13

OpenStudy (ybarrap):

Perfect! Now let's all three results and see what we get...

OpenStudy (ybarrap):

Now let's add all three results and see what we get ...

OpenStudy (ybarrap):

$$ \cfrac{6}{14}\times\cfrac{5}{13}+\cfrac{8}{14}\times\cfrac{6}{13}+\cfrac{6}{14}\times\cfrac{8}{13}=? $$

OpenStudy (leozap1):

9/13 (So 69%?)

OpenStudy (ybarrap):

Yes!

OpenStudy (ybarrap):

Here is the easier way...

OpenStudy (ybarrap):

We know that P(gg) + P(gb) + P(bg) + P(bb) = 1 So P(gg) + P(gb) + P(bg) = 1- P(bb) This means we really just needed to compute P(bb). Do you see that?

OpenStudy (ybarrap):

$$ P(gg) + P(gb) + P(bg) = 1- P(bb)=1-\cfrac{8}{14}\times\cfrac{7}{13} $$

OpenStudy (ybarrap):

And this equals 9/13. Much easier approach

OpenStudy (leozap1):

Okay thank you! That makes a lot more sense, and will make it easier to solve future problems. :)

OpenStudy (ybarrap):

You're welcome